OpenStudy (anonymous):

Simple Interest Question? I'm stuck. Arlene Frank is saving money for her college education. She deposited some money in a savings account paying 5% and $1200 less than that amount in a second account, paying 4%. The two accounts produced a total of $141 interest in 1 year. How much did she incest at each rate? -I have tried Principal * rate, and I keep ending up with a mess with my problem set up. Can someone explain all of this to me? Thanks.

OpenStudy (anonymous):

Invest at each rate* Sorry. Typo.

OpenStudy (anonymous):

try to google search-step by step then what you need to know

OpenStudy (anonymous):

I have tried google. Nothing shows up

OpenStudy (anonymous):

0.05(X) + 0.04(X-1500)=141 solve for X d thus X-1500 0.09 X = 141 + 60 = 201 X = 2233 and x - 1500= 733. Check 0.05(2233) + 0.04(733) = 141 Yes.

OpenStudy (anonymous):

First line should have printed as 0.05 X + 0.04(X-1500) = 141

OpenStudy (anonymous):

Where does the 1500 come from?

OpenStudy (anonymous):

It says in the book that the answer is $2100 at 5%, and $900 at 4%????

OpenStudy (anonymous):

My error. I should have used 1200, not 1500. I misread it. Sorry. Replace my 1500 with 1200 and you will get the right answer.